4-5: Profitability Ratios BEP, ROE, and ROIC Duval Manufacturing recently reported the following information ROA Interest expense Accounts payable $655,000 9% $196,500 $950,000 Duvals tax rate is 30%. Duval finances with only debt and common equity, so it has no preferred stock, 40% of its total invested capital is debt, while 60% of its total invested capital is common equity. Calculate its basic earning power (BEP), its return on equity (ROE), and its return on invested capital (ROIC) Round your answers to two decimal places. ROE ROIC
Answer #1

1.        Basic Earning Power (BEP)

           It is the ratio business’s income before tax and financial leverage to Total Assets.

          

           BEP = EBIT/Total Assets

2.        Return on Equity (ROE)

It is ratio between company’s income available to shareholders to average equity capital

          

           ROE = Net Income/Equity capital

3.        Return on Invested capital (ROIC)

          

It is the ratio between Earning before interest and tax (EBIT) to total capital employed.

                        ROIC = EBIT/(Debt+Equity)

Now, It is assumed that provided Net Income $ 655,000 is income available to shareholders which means it is Earning after interest and tax.

Following Information provided -

Net Income - $ 655,000

Return on Assets(ROA) - 9%

Interest Expenses (I) = $ 196,500

Accounts payable = $ 950,000

Tax Rate = 30%

ROA = Net Income/Total Assets

0.09 = 655,000/Total Assets

Total Assets = 655,000/0.09

Total Assets = $ 7,277,777.78

EBIT = Net Income/(1-tax) + Interest expenses

= 655,000/(1-0.3) + 196,500

= $ 1,132,214.29

1.

BEP = EBIT/Total Assets

= 1,132,214.29/7,277,777.78

= 0.15557

=15.56%

Thus, BEP is 15.56%.

2.

Debt + Equity = Total Assets - Accounts Payable

Debt + Equity = 7,277,777.78 - 950,000

= 6,327,777.78

Debt is 40% of invested capital and Equity is 60% of invested capital.

Thus, Equity = 6,327,777.78*0.6

= 3,796,666.67

Debt = 6,327,777.78*0.4

= 2,531,111.11

ROE = Net Income/Equity

= 655,000/3,796,666.67

= 0.1725

= 17.25%

Thus, Return on Equity is 17.25%

3.

Invested Capital (Capital employed) = Debt + Equity

= 6,327,777.78

ROIC = EBIT/Capital Employed

= 1,132,214.29/6,327,777.78

= 0.1789

= 17.89%

Thus, Return on Invested capital is 17.89%.

Please note some analyst ignore cost of borrowing while calculating return on Assets (ROA), In such case -

ROA = (Net Income + Interest)/Total Assets